### v1

# SeniorTalk: A Chinese Conversation Dataset with Rich Annotations for Super-Aged Seniors

Abstract
While voice technologies increasingly serve aging populations, current systems exhibit significant performance gaps due to inadequate training data capturing elderly-specific vocal characteristics like presbyphonia and dialectal variations. The limited data available on super-aged individuals in existing elderly speech datasets, coupled with overly simple recording styles and annotation dimensions, exacerbates this issue. To address the critical scarcity of speech data from individuals aged 75 and above, we introduce SeniorTalk, a carefully annotated Chinese spoken dialogue dataset. This dataset contains 55.53 hours of speech from 101 natural conversations involving 202 participants, ensuring a strategic balance across gender, region, and age. Through detailed annotation across multiple dimensions, it can support a wide range of speech tasks. We perform extensive experiments on speaker verification, speaker diarization, speech recognition, and speech editing tasks, offering crucial insights for the development of speech technologies targeting this age group.
SeniorTalk: A Chinese Conversation Dataset with Rich Annotations for Super-Aged Seniors
Yang Chen1+, Hui Wang1+, Shiyao Wang1, Junyang Chen1,Jiabei He1, Jiaming Zhou1, Xi Yang2, Yequan Wang2, Yonghua Lin2, Yong Qin1thanks: Yong Qin is the corresponding author. 1College of Computer Science, Nankai University, 2Beijing Academy of Artificial Intelligence, Beijing, China , Correspondence: 2120230617@mail.nankai.edu.cn, qinyong@nankai.edu.cn
First Author and Second Author contribute equally to this work.
1 Introduction
Dataset Age Style Annotation Features
Speaker Region Transcript Timestamp Accent Level Sound Event
ASR0060 55+ Reading North/South/Other $\checkmark$ N/A $\times$ $\times$
MECSD 55+ Reading $\times$ $\checkmark$ N/A $\times$ $\checkmark$
SeniorTalk 75+ Conversation Provincial $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 1: Comparison of Chinese elderly speech datasets, including AISHELL-ASR0060 (marked in the table as ASR0060), MECSD, and SeniorTalk, across annotation features.
The rapid global aging population presents both significant challenges and opportunities in the development of technologies specifically designed for older adults, particularly those aged 75 and over (Scuteri and Nilsson, 2024; Chumuang et al., 2024). As the number of people in this ultra-high-age group continues to grow, it becomes increasingly important to enhance the accessibility and inclusion of speech-based technologies for this demographic (Young and Mihailidis, 2010; Portet et al., 2013). However, many state-of-the-art speech systems struggle to perform effectively within the elderly population, exhibiting biases on elderly vocal patterns. Kulkarni et al. (2024); Feng et al. (2021). For instance, existing speech recognition systems often show poor performance with older users, influenced by factors such as speech deterioration, hearing loss, health issues (Fraser et al., 2015), and the diversity of speech patterns among older adults (Geng et al., 2022; Hu et al., 2024). A key underlying reason for this challenge is the lack of datasets specifically addressing the unique needs of ultra-high-age individuals Hu et al. (2024), which hampers the development of robust foundation models and the application of tailored solutions.
Although existing studies have made some efforts to collect geriatric speech data (Sekerina et al., 2024; Fukuda et al., 2020), significant limitations remain. First, current speech corpora predominantly focus on younger adults or healthy senior populations, with samples that consist mainly of scripted speech and exhibit standardized accents. For example, both the AISHELL-ASR0060 and MECSD datasets (Wang et al., 2019) adopt an unusually low age threshold of 55 years for defining elderly participants. Similarly, the S-JNAS corpus of elderly Japanese speech reports a mean speaker age of 67.6 years (Fukuda et al., 2020). These fall below the World Health Organization’s (WHO) geriatric classification designating 75+ years as the late elderly phase which is a period associated with progressive age-related decline in physiological function (Orimo et al., 2006).
Second, the current collection paradigms and annotation methods of existing datasets further limit their practical applicability. A large portion of speech resources for the elderly primarily focuses on reading style (Wang et al., 2019; Iribe et al., 2015), which do not reflect the everyday communication scenarios that elderly people encounter in real life. Furthermore, many corpora are tailored for narrowly defined tasks such as automatic speech recognition (ASR) or pathology detection (Wang et al., 2019), thereby preventing comprehensive characterization of age-related vocal variations. These datasets also lack key features such as speaker diarization or dialect labels, which restrict their ability to support a wider array of use cases and hinder their robustness in addressing the diverse challenges of speech processing.
To address these limitations, we introduce SeniorTalk, a Mandarin dataset consisting of spontaneous conversations among individuals aged 75 and older. As shown in Table 1, this dataset comprises natural conversational recordings from 202 native Chinese speakers, representing a rich diversity of regional, age, and gender demographics, and captured in authentic, real-world interaction settings. It effectively addresses the current gaps and limitations in datasets focused on elderly populations, particularly the underrepresentation of super-aged seniors and the lack of diversity in recording styles and annotation dimensions. Moreover, we conduct extensive experiments across various speech tasks, providing a benchmark specifically for the elderly population. By open-sourcing this corpus along with fine-grained metadata, we aim to bridge the vocal age gap and promote the development of equitable voice technologies for aging societies. SeniorTalk makes three key contributions to the field:
  • We recruit 202 super-elderly speakers from 16 provinces in China, ensuring balance across gender, age, and geography. This resulted in SeniorTalk, a dataset with 55.53 hours of data from 101 conversations.
  • We provide detailed, multi-dimensional annotations for the dataset, encompassing speaker information, transcriptions, timestamps, and more. These annotations enable comprehensive speech signal analysis and support a wide range of speech-related tasks for the elderly.
  • We conduct a comprehensive series of experiments using the dataset, covering tasks such as speaker verification, speaker diarization, speech recognition, and others. This establishes a solid benchmark for evaluating models across various speech-related tasks.
2 Related Work
Corpus Language Age Style # Spks. Dur.(hrs) Year Avail.
BraPoRus (Sekerina et al., 2024) Brazilian Portuguese-Russian 59-98 Monologue, interview, … 1,500 170 2024 N
EARS (Fukuda et al., 2023) Japanese 70-99 Reading 123 13.4 2023 N
Improving S-JNAS (Fukuda et al., 2020) Japanese 65-99 Reading 221 31.7 2020 Y
AISHELL-ASR0060 Mandarin 55+ Reading 503 793 2019 Y
MECSD (Wang et al., 2019) Mandarin 55-85 Reading 85 110 2019 P
elderLUCID (Hazan et al., 2017) English 19-84 Reading 83 N 2017 N
Develop S-JNAS (Iribe et al., 2015) Japanese 60-98 Reading 100 9.2 2015 Y
ERES38 (Aman et al., 2013) French 68-98 Interview 22 17 2013 N
AD80 (Aman et al., 2013) French 62-94 - 43 4.7 2013 N
CCC (Pope and Davis, 2011) English 65+ Interview 600+ 800+ 2011 Y
S-JNAS Japanese 60-90 Reading 301 - 2007 Y
E-MIC Korean 65-85 conversation 100 3 - Y
Table 2: Summary of related elderly speech datasets. Key characteristics include the age ranges of speakers (Age), the number of speakers (# Spks.), publication year (Year), and availability status (Avail.), where ’P’ indicates partial availability.
Several corpora have been developed to address the specific acoustic characteristics of elderly speech. Early work in this area includes the Japanese Newspaper Article Sentences Read Speech Corpus of the Aged (S-JNAS), a foundational resource for Japanese elderly speech research. Subsequent research has broadened the scope of investigation, with efforts such as the Carolinas Conversations Collection (CCC) (Pope and Davis, 2011) focusing on multiethnic elderly speakers with chronic conditions, offering valuable insights into how sociolinguistic factors and health status influence speech production. The AD80 and ERES38 corpora (Aman et al., 2013) advance French elderly speech analysis through distress detection benchmarks for ambient assisted living systems. Further corpus development has continued with the creation of specialized datasets, including a corpus of 100 elderly Japanese speakers designed to enhance human-robot interaction in elder care (Iribe et al., 2015). The elderLUCID project (Hazan et al., 2017) examines the complexities of speech communication in older adults, considering the interplay of hearing loss, phonation, articulation, and cognitive factors.
More recently, researchers have focused on specific languages and conditions, as exemplified by the Mandarin Speech Database for Early Dementia Detection (MECSD) (Wang et al., 2019) and the AISHELL-ASR0060 databasehttps://www.aishelltech.com/General_Datasets for elderly Mandarin speech. Improvements to existing corpora, such as S-JNAS, have also been explored, including the creation of acoustic models specifically designed for "super-elderly" speakers (Fukuda et al., 2020). The Elderly Multimodal Interpersonal Conversation (E-MIC)https://ai4robot.github.io/emic-en/ dataset expands the scope of analysis by incorporating multimodal data, including video and audio, to study turn-taking in elderly conversations. Further work, such as EARS (Fukuda et al., 2023), has continued to refine acoustic modeling techniques for super-elderly Japanese speakers, while the BraPoRus corpus (Sekerina et al., 2024) highlights the importance of preserving heritage languages and the challenges of remote data collection during the COVID-19 pandemic.
3 Dataset description
3.1 Dataset overview
SeniorTalk is designed specifically for the ultra-elderly population aged 75 and above, offering a comprehensive collection of spoken dialogue data aimed at supporting various speech-related tasks. The dataset includes a total of 101 recorded speech dialogues, representing a diverse range of linguistic characteristics. It spans 55.53 hours of speech data, recorded from 202 participants, and features 60,029 individual utterances. Additionally, the dataset is enriched with annotations across 8 distinct dimensions, ensuring its suitability for training and evaluating robust models across multiple speech-processing tasks.
3.2 Statistics
Figure 1: A stacked bar chart showing the population count by age and gender. The chart illustrates the distribution of male and female populations across different age groups.
Figure 2: A dual-layer pie chart representing the distribution of provinces across the North and South regions. The inner layer highlights the overall proportions of North and South, while the outer layer shows the specific distribution of provinces within each region.
Participants
We recruit 202 participants aged between 75 and 85 years, ensuring a diverse representation across different segments of the ultra-elderly population, and obtain the necessary consent for their participation. Specific authorization details are provided in Appendix A.1. Figure 1 illustrates the gender distribution across different age groups. The dataset includes 69 male and 131 female speakers, with a higher proportion of females. This gender imbalance stems from the relative ease of recruiting female participants during the data collection process, likely due to the higher life expectancy of women in the targeted age range. Geographically, as shown in Figure 2, 94 participants are from northern China, while 108 are from southern China, covering regions such as Beijing, Shanghai, and Sichuan. This regional diversity enhances dialect recognition models by exposing them to a wide range of linguistic patterns.
Figure 3: This histogram visualizes the frequency distribution of the given dataset, showing how the values are spread across different ranges.
Recording
The dataset comprises spontaneous speech dialogues that span a wide range of topics, aimed at capturing the natural flow and diversity of real-life conversations. These dialogues are recorded using various mobile devices, with a distribution of 70% from Android devices and 30% from iOS smartphones. The topics of these dialogues are specifically chosen to address relevant issues for older adults, such as health, pets, retirement, and other related matters. Each conversation typically covers between one and three topics, with the distribution of topics visualized in Appendix A.2. Figure 3 illustrates the distribution of session durations, which range from 25 to 50 minutes, with the majority of sessions clustering around the 35-minute mark.
Annotation
The dataset is annotated by a team of 14 trained annotators who undergo unified training. During the training, they are instructed on a well-defined annotation workflow and strict standards to ensure consistency and accuracy. Detailed guidelines on the annotation process and standards are provided in Appendix A.3.
Annotations are made across four main dimensions. At the speaker profile level, attributes such as age, gender, and origin are annotated. These annotations are particularly useful for analyzing elderly speech signals, helping to study age-related speech characteristics. The session level includes annotations for temporal segmentation and overlapping speech, which are essential for segmenting speech and identifying overlaps in multi-speaker scenarios. At the utterance level, transcriptions and accent intensity are annotated. Transcriptions support ASR. Finally, at the token level, special sound events like laughter are marked, providing insight into non-verbal communication.
4 Experiments
In this section, we assess our dataset across various tasks, including speaker verification, speaker diarization, speech recognition, and speech editing.
Model # Params Dim Dev (%) PLDA Cosine similarity
EER (%) minDCF EER (%) minDCF
x-vector 4.2M 512 12.04 14.63 0.9768 19.26 0.9598
ResNet-TDNN 15.5M 256 4.372 10.88 0.8450 11.50 0.9196
ECAPA-TDNN 20.8M 192 8.86 11.54 0.9214 10.24 0.9582
Table 3: Results of fine-tuning baselines on the speaker verification task, where Dim indicates the dimension of the extracted embeddings and Dev represents the EER on the validation set.
4.1 Speaker Verification
This section introduces the Speaker Verification (SV) task, which is essential for verifying the identity of speakers within the geriatric population to ensure their financial security. To facilitate this task, we implement a data partitioning strategy, with further details provided in Appendix B.1.1.
4.1.1 Metrics
We adopt two scoring approaches: probabilistic linear discriminant analysis (PLDA) (Prince and Elder, 2007) and cosine similarity, with evaluation based on two metrics: (1) Equal Error Rate (EER): We define a threshold $\tau$ where the miss probability equals the false alarm probability. Specifically, if the similarity score is above this threshold, the system accepts that the speakers are the same person; if it is below this threshold, the system rejects the claim. This threshold is selected when the false acceptance rate equals the false rejection rate. (2) Minimum Detection Cost (minDCF): A cost-sensitive metric for evaluating speaker verification systems under application-specific conditions.
4.1.2 Baselines
We adapt three state-of-the-art speaker embedding systems pre-trained on VoxCeleb (Nagrani et al., 2017) through domain-specific fine-tuning via SpeechBrain (Ravanelli et al., 2021): x-vector architecture (Snyder et al., 2018)https://huggingface.co/speechbrain/spkrec-xvect-voxceleb, ECAPA-TDNN (Desplanques et al., 2020)https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb, and ResNet-TDNN (Villalba et al., 2020)https://huggingface.co/speechbrain/spkrec-resnet-voxceleb. Detailed hyperparameters are listed in Appendix B.2.3.
4.1.3 Results and Analysis
Table 3 demonstrates two critical findings from our geriatric voice analysis: First, the strong baseline performance confirms our dataset’s suitability for aging voice biometrics. However, age-related vocal degradation (e.g., pitch instability, articulatory imprecision) introduces distinct challenges compared to pediatric voices, potentially affecting gender differentiation and speaker discriminability.
Second, the ECAPA-TDNN’s relative underperformance versus parameter-efficient model ResNet using PLDA testing method suggests susceptibility to overfitting in limited elderly speech data scenarios. This emphasizes the need for data augmentation strategies, regularization techniques and hyperparameter finetuning when deploying deep speaker models in geriatric voice applications.
Model # Params Dim collar=0 collar=0.25
DER (%) Confusion (%) DER (%) Confusion (%)
ResNet-34-LM 15.5M 256 33.14 16.82 28.39 16.85
x-vector 4.2M 512 53.01 36.69 49.82 38.28
ResNet-TDNN 15.5M 256 43.44 27.13 39.58 28.03
ECAPA-TDNN 20.8M 192 27.84 11.52 22.85 11.31
Table 4: Results of fine-tuning speaker embedding extraction models and pretrained model ResNet-34-LM on the speaker diarization task.
4.2 Speaker Diarization
This section introduces the speaker diarization task, which entails partitioning audio recordings into segments that correspond to individual speakers. This task is essential for assessing the performance of various speaker models, as detailed in Section 4.2.3. The data split is the same as speaker verification task.
4.2.1 Metrics
We employ the Diarization Error Rate (DER) as the evaluation metric for the speaker diarization task. The computation of DER is defined as follows:
${DER=\frac{FA+MD+Conf}{T},}$ (1)
where $FA$ denotes the number of segments incorrectly identified as speech when no speaker is present, $MD$ represents the number of segments where speech is present but not detected. $Conf$ refers to the number of segments where detected speech is attributed to the wrong speaker, and $T$ indicates the total number of speech segments in the reference transcript.
In Table 4, we explore two collar settings: 0 seconds and 0.25 seconds. The collar parameter defines a time window around the detected speaker boundaries, allowing for a margin of error in segment alignment. A collar value of 0 seconds requires exact matching of boundaries, while a value of 0.25 seconds introduces a quarter-second tolerance to accommodate minor discrepancies in detection.
4.2.2 Baselines
For the speaker diarization task, we use the PyAnnote toolkit (Bredin, 2023; Plaquet and Bredin, 2023) https://github.com/pyannote/pyannote-audio. This speaker diarization pipeline consists of three primary components: Voice Activity Detection (VAD), the speaker extractor, and clustering methods (e.g., K-Nearest Neighbors).
  • VAD: We employ the pyannote/segmentation-3.0 model https://huggingface.co/pyannote/segmentation-3.0, an end-to-end neural architecture for joint speech activity detection and speaker segmentation. Since false alarm and missed detection rates remain consistent across experimental conditions due to fixed VAD parameters, Table 4 exclusively reports confusion errors and DER metrics.
  • Speaker extractor: For the speaker extractor module, we replace the default ResNet-34-LM extractor from PyAnnote https://huggingface.co/pyannote/wespeaker-voxceleb-resnet34-LM with fine-tuned versions of x-vector, ECAPA-TDNN, and ResNet-TDNN architectures, all adapted for speaker verification. This yields four distinct experimental configurations, as detailed in Table 4.
  • Clustering method: For the clustering method, we employ the default Spectral Clustering, as proposed by PyAnnote.
4.2.3 Results and Analysis
Our experimental analysis yields three principal conclusions regarding elderly speaker diarization:
First, after fine-tuning with the elderly dataset, the ECAPA-TDNN model outperforms the PyAnnote default model ResNet-34-LM by 5.3% DER improvement at 0 collar and 5.54% DER imporvement at 0.25 collar. This superiority demonstrates the effectiveness of our dataset for the diarization of elderly speakers.
Second, our empirical analysis on speaker diarization reveals that ECAPA-TDNN significantly outperforms both ResNet-TDNN and x-vector-based systems. However, Table 3 demonstrates that these performance gaps diminish in speaker verification tasks, with ResNet-TDNN even surpassing ECAPA-TDNN under PLDA scoring. We attribute this discrepancy to two primary factors: the shift to a conversational dataset and ECAPA-TDNN’s inherent architectural advantages in noise resilience. While ECAPA-TDNN is typically trained on clean speech corpora, our experiments evaluate its robustness in realistic scenarios characterized by background noise and overlapping speech. The model’s noise-resilient components enhance its feature extraction capabilities under such adverse conditions, which proves critical in diarization tasks but less impactful in controlled speaker verification settings. This contrast highlights the importance of conversational datasets in assessing model robustness, as they better reflect real-world challenges that differentiate model performance.
Third, our experiments reveal notably high DER and confusion errors. Our analysis attributes this phenomenon to two dataset-specific factors: (1) a pronounced gender imbalance (1:3 female-to-male ratio) contrasting standard benchmarks’ balanced distributions, and (2) age-related vocal changes in elderly speakers that reduce the saliency of secondary sexual voice characteristics. These combined effects create systematic challenges for speaker identity separation, particularly in conversational contexts where demographic diversity and physiological aging patterns naturally occur.
4.3 Speech Recogition
Encoder # Params CER Accent Region
No Light Moderate Heavy South North
Transformer 14.1M 48.99 22.58 49.05 51.07 80.95 48.5 50.24
Conformer 15.7M 34.61 21.23 34.21 37.62 59.52 34.55 34.74
E-Branchformer 16.9M 33.25 20.71 33.03 35.32 64.29 32.97 33.94
Table 5: Decoding performance (CER, %) of Transformer, Conformer, and E-Branchformer models using Attention rescoring, with accent differentiation and region categorization.
Model # Params Zero-shot Fine-tuning
Paraformer-large 232M 14.91 14.41
Whisper-tiny 39M 92.20 58.80
Whisper-base 74M 64.02 38.17
Whisper-small 244M 55.83 28.69
Whisper-medium 769M 60.47 25.77
Whisper-large-v3 1,550M 57.74 23.84
Table 6: Character Error Rate (CER) (%) of supervised pre-trained baselines in both zero-shot and fine-tuning settings.
Automatic speech recognition entails transcribing spoken language into text, and recognizing elderly speech patterns is crucial in emergency response scenarios due to age-related vocal characteristics that can affect system reliability. This section presents an empirical evaluation of the collected elderly speech corpus, with data partitioning strategies detailed in Appendix B.1.2 .
4.3.1 Metrics
The experimental results demonstrate the performance on the test dataset after training with the train dataset, using Character Error Rate (CER) as the evaluation metric, which is computed by the following equation:
$CER=\frac{S+D+I}{N},$ (2)
where S, D, and I respectively signify the quantities of substitutions, deletions, and insertions. denotes the cumulative number of characters within the reference text. When evaluating character-level transcription accuracy, a system featuring a lower CER is typically regarded as more proficient.
4.3.2 Baselines
We employ the open-source wenet toolkit (Yao et al., 2021) as our training framework, selecting Transformer, Conformer, and E-Branchformer as our baselines. All models are trained using a combined Connectionist Temporal Classification (CTC) and Attention-based Encoder-Decoder (AED) approach. We fine-tune the hyperparameters of these three models to ensure comparable parameter counts while achieving optimal performance. Detailed hyperparameter configurations can be found in the Appendix B.2.1 and B.2.2.
The following models are considered in the context of training from scratch:
  • Transformer: The standard Transformer architecture employs both CTC and AED objectives, establishing a widely-adopted baseline for ASR.
  • Conformer: The Conformer (Gulati et al., 2020) model integrates convolutions with self-attention for ASR, sandwiched between two feed-forward layers.
  • E-Branchformer: Proposed by Kwangyoun Kim et al. (Kim et al., 2022), E-Branchformer builds upon the Branchformer (Peng et al., 2022), which attains performance levels comparable to Conformer. E-Branchformer improves on this framework by implementing a novel merging strategy and integrating additional point-wise modules.
In addition to the three models trained from scratch, we fine-tune two pre-trained models: Paraformer, which employs the Wenet framework, and Whisper, whose hyperparameters and code base are described in Appendix B.2.2.
  • Paraformer: Proposed by Gao et al. (Gao et al., 2022), Paraformer is a fast and accurate parallel transformer model that leverages a continuous integrate-and-fire (CIF) (Dong and Xu, 2020) predictor to estimate the number of tokens and generate hidden representations. This pre-trained model is trained on a non-public, industry-grade dataset comprising 60,000 hours of Chinese ASR data.
  • Whisper: Whisper (Radford et al., 2023) is a Transformer-based multilingual ASR model developed by OpenAI, trained on 68,000 hours of labeled speech data. We examine various versions of Whisper, ranging from tiny to large-v2, with model sizes varying from 39M to 1.55Bhttps://github.com/openai/whisper.
4.3.3 Results and Analysis
Models Trained from Scratch
We analyze models trained from scratch across three aspects: baseline performance, accent intensity impact, and regional variations.
Baseline Model Performance: Table 5 compares three models, namely E-Branchformer, Conformer, and Transformer, which are trained from scratch and utilize the attention rescoring decoding method. The E-Branchformer achieves the lowest overall CER, outperforming Conformer by 1.36% and Transformer by 15.74%. This performance advantage holds consistently across all tested conditions, including varying accent intensities and geographical regions.
Accent Intensity Impact: As shown in Table 5, CER increases with accent intensity: Moderate Accent yields higher CER than Light Accent, which in turn exceeds No Accent. This trend highlights the growing recognition challenge as accents become more pronounced.
Regional Variations: The E-Branchformer achieves a 0.97% lower CER in the South than in the North, with the Conformer and Transformer showing improvements of 0.19% and 1.74%, respectively. These differences indicate slightly higher recognition difficulty in the North, although overall performance remains comparable. At the provincial level, substantial variations in CER are observed across different provinces. For detailed analysis, please refer to the AppendixB.3.
Finetuned Models
Table 6 presents comparative CER results for two model families: 1) our Wenet-finetuned Paraformer-large architecture (Yao et al., 2021), and 2) a series of Whisper models (Radford et al., 2023) adapted through whisper-flamingo fine-tuninghttps://github.com/roudimit/whisper-flamingo. The analysis reveals two key findings:
First, although Whisper models show modest CER improvements with larger parameter counts (except for small), all variants have CERs exceeding 50% in zero-shot senario. This performance gap suggests potential domain mismatch between Whisper’s training data distribution and elderly speech characteristics, possibly due to underrepresentation of senior voices in Whisper’s pretraining corpus. Such mismatch may explain the frequently observed hallucination patterns in elderly speech recognition results. Notably, our targeted fine-tuning reduces CERs substantially, improving cer of Whisper-large-v3 from 57.74 to 23.84%, demonstrating both the challenge level of our elderly speech dataset and its utility for domain adaptation.
Second, the Paraformer-large model achieves 14.91% CER in zero-shot evaluation and 14.41% CER after fine-tuning , outperforming all Whisper variants by significant margins. This advantage likely stems from Paraformer’s pretraining on 60,000 hours of proprietary Chinese speech data encompassing diverse regional accents, age groups, and speaking styles. This demonstrates that Paraformer has better generalization ability in the field of Chinese speech recognition.
Method MCD (↓) STOI (↑) PESQ (↑)
CampNet 7.302 0.220 1.291
EditSpeech 6.225 0.514 1.363
A3T 5.851 0.586 1.455
FluentSpeech 5.811 0.627 1.645
Table 7: Evaluation results of speech editing models trained from scratch.
4.4 Speech Editing
Speech editing is a generative task that modifies corresponding speech based on text alterations, and editing speech from elderly individuals is particularly useful for editing interviews with seniors. To facilitate this task, we implement a new data split, as detailed in Appendix B.1.3. However, utilizing this data for generative tasks poses risks related to elder fraud, particularly in the context of telemarketing scams targeting senior individuals.
4.4.1 Baselines
We employ the Speech Editing Toolkit https://github.com/pyannote/pyannote-audio framework to implement our models, which include CampNet(Wang et al., 2022), EditSpeech(Tan et al., 2021), A3T(Bai et al., 2022) , and FluentSpeech(Jiang et al., 2023). The detailed hyperparams are described in Appendix B.2.4.
4.4.2 Metrics
We utilize Mel-Cepstral Distortion (MCD) Kubichek (1993), Short-Time Objective Intelligibility (STOI) Taal et al. (2010), and Perceptual Evaluation of Speech Quality (PESQ) Rix et al. (2001), which are commonly used objective metrics for assessing speech generation quality.
4.4.3 Performance Analysis
The results presented in Table 7 indicate that FluentSpeech achieved the best performance, with all three objective metrics falling within acceptable ranges. In relevant research (Jiang et al., 2023; Bai et al., 2022; Wang et al., 2022; Liu et al., 2024), these experimental values are widely recognized as acceptable. This suggests that our dataset is suitable for generative tasks.
5 Conclusion
In this paper, we present SeniorTalk, a Mandarin dataset featuring spontaneous conversations among individuals aged 75 and older. With 55.53 hours of data from 202 speakers across 16 provinces in China, this dataset offers valuable resources for developing voice technologies for aging populations. By providing detailed annotations and conducting extensive experiments across key speech tasks, we establish SeniorTalk as a benchmark for evaluating models for elderly speakers, aiming to bridge the vocal age gap and promote the development of more inclusive voice technologies.
Limitations
While SeniorTalk advances research on elderly speech processing, several limitations should be noted. First, while our dataset includes 202 participants across 16 provinces, its scale remains modest compared to the linguistic diversity and population size of China. Expanding the dataset to include thousands of speakers would better capture regional accents, dialects, and individual vocal variability. Second, the current age range could be extended to include more super-elderly individuals (e.g. 85+ or centenarians), as advanced aging often correlates with unique vocal patterns due to physiological decline, cognitive changes, or age-related pathologies. Third, geographic coverage could be broadened to encompass underrepresented regions (e.g., rural areas or ethnic minority regions) to mitigate potential demographic biases. Additionally, longitudinal data capturing vocal aging trajectories could further enhance the dataset’s utility. These extensions would strengthen the generalizability of voice technologies for aging populations and address the complex interplay of age, health, and regional linguistic diversity.
Ethics Statement
This study strictly adhered to rigorous ethical protocols to safeguard elderly participants’ rights. Audio recordings were conducted in quiet indoor environments at senior care facilities. To accommodate potential cognitive decline among participants, multiple conversation topics were provided during dyadic interactions, with recording devices positioned equidistant between paired participants.
Prior to each session, informed consent was obtained after explaining the research objectives and data collection parameters, including voice characteristics, conversational content, age documentation, and accent analysis. Participants were compensated monetarily, with amounts ranging from 330 to 400 RMB. This compensation was calibrated according to local purchasing power, thereby ensuring equitable remuneration tailored to the specific geographical locations of the participants. All personal identifiers (e.g., national ID numbers, full names) were systematically de-identified by replacing them with unique speaker identifiers, despite initial age verification requiring temporary ID inspection.
The dataset carries inherent risks requiring stringent governance. Potential malicious exploitation for voice synthesis could potentially exacerbate elderly-targeted telecommunications fraud through voice spoofing. Accordingly, access is strictly restricted to vetted academic researchers through institutional credential verification, with legally binding agreements prohibiting commercial use or redistribution.
Key ethical safeguards implemented include: 1) Explicit participant consent through verbal and written confirmation, 2) Comprehensive privacy preservation via anonymization protocols, 3) Fair compensation aligned with regional economic standards, 4) Institutional Review Board approval for all data collection procedures, 5) Multi-layered access controls preventing unauthorized usage.
This framework ensures compliance with international research ethics standards (Declaration of Helsinki, GDPR principles) while balancing scientific utility with participant protection in vulnerable populations.
Appendix A Supplementary Dataset Information
A.1 Authorization
This study employs time-series biometric data collected under a formal ethical authorization framework developed in collaboration with a third-party AI data service provider specializing in biometric acquisition. The framework ensures methodological transparency and regulatory alignment through the following mechanisms:
Legally Binding Consent Protocols Participants provide informed consent for the collection of vocal and physiological time-series data, explicitly authorizing its use in AI research and derivative model development.
Rights Management All datasets and derived models remain the exclusive intellectual property of the anonymized provider. Participants retain conditional rights to access, modify, or request data deletion, contingent on technical feasibility (deletion requests that invalidate associated research outputs may require proportional compensation).
Cross-Jurisdictional Compliance Third-party data sharing requires explicit opt-in consent, with passive approval mechanisms activated only after a 72-hour objection period following notification.
Biometric Corpus Specifications The dataset includes anonymized temporal features such as age, gender, regional dialect markers, and sequential vocal patterns (e.g., pitch dynamics, spectral entropy trajectories).
A.2 Topic
This section presents the frequency statistics of all topics in the dataset, as illustrated in Figure 4. In total, there are 13 major categories, encompassing 58 distinct topic labels. The distribution of these 13 categories is depicted in the figure. These topics primarily reflect the everyday concerns of elderly individuals, with particular emphasis on areas such as Leisure, Health, and Retirement Life, which appear most frequently. These topics align closely with the key interests and priorities of older adults. Understanding and focusing on these subjects is crucial, as it can greatly benefit the performance of ASR systems, especially when catering to elderly users in relevant contexts.
A.3 Annotation
This section presents information facilitating annotation, covering details about annotators, the annotation process, and the content being annotated.
ID Region Age Gender Education
1 Henan 36 Male Bachelor
2 Hebei 23 Female Bachelor
3 Hebei 24 Female Bachelor
4 Hebei 24 Female Bachelor
5 Hebei 30 Female Bachelor
6 Fujian 27 Male Bachelor
7 Fujian 23 Male Bachelor
8 Fujian 23 Female Bachelor
9 Fujian 38 Female Bachelor
10 Chuzhou 22 Male Bachelor
11 Chuzhou 27 Male Bachelor
12 Chuzhou 23 Female Bachelor
13 Yunnan 27 Male Bachelor
14 Hunan 47 Male Bachelor
Table 10: Annotator Information
A.3.1 Annotator
As shown in Table 10, We have a total of 14 annotators. Eleven of them are between 23 and 30 years old, and three are between 30 and 50 years old. All annotators hold undergraduate degrees. Moreover, they are proficient in their local dialects, which facilitates the transcription of accented speech.
A.3.2 Annotation workflow
The annotation process utilizes a proprietary cloud-based platform developed by a third-party data service provider. Annotators access tasks through this platform, which facilitates manual segmentation of conversational data into discrete utterances followed by automated speech recognition to generate preliminary transcripts. These baseline outputs undergo human revision alongside annotation of linguistic features including accent intensity levels (see Table 8) and paralinguistic markers such as [NOISE] or [MUSIC]. A multi-tier quality assurance protocol is implemented through the same platform, where dedicated reviewers systematically validate annotation consistency and accuracy prior to dataset finalization.
A.3.3 Annotation Information
Table 9 presents our four-tier annotation framework comprising speaker, session, utterance, and token levels, with each layer supporting distinct speech processing tasks through systematically designed metadata.
The speaker-level annotations capture demographic attributes (geographic origin, gender, age) to enable geriatric voice analysis.
At the session level, comprehensive temporal annotations of turn-taking boundaries and overlap detection facilitate speaker diarization and speech separation benchmarking.
Utterance-level transcriptions incorporate accent intensity labels and orthographic normalization, jointly supporting automatic speech recognition and ordinal classification of vocal aging patterns.
The token tier extends this granularity with paralinguistic markers, including [MUSIC], [NOISE], [LAUGHTER], and [SONANT], enabling quantitative analysis of non-lexical vocal characteristics critical for elderly communication studies.
Appendix B Supplementary Experiments
This section mainly presents supplementary experimental details, including the dataset partitioning for each task and the hyperparameters. The fine-tuning and zero-shot tests of the Whisper model are performed using an NVIDIA A800, while all other experiments are carried out with an NVIDIA GeForce RTX 3090.
B.1 Data Split
Split # Spk. # Utt. Dur. (hrs) Avg. (s)
Train 0-182 48591 30.47 2.26
Dev 0-182 5398 3.40 2.27
Test 182-202 6040 3.95 2.35
Sum 0-202 60029 37.82 2.27
Table 11: Summary of dataset splits, including the speakers range (# Spk.), utterances (# Utt.), total duration (Dur.), and average utterance length (Avg.) for speaker verification and speaker dirazation tasks.
Encoder Accum Grad Batch size LR Warmup Epochs
Transformer 4 4 1.00E-03 5000 100
Conformer 4 4 1.00E-03 5000 60
E-Branchformer 4 4 5.00E-04 5000 60
Table 12: Hyperparameters for training ASR models from scratch.
Model Accum Grad Batch size Learning rate Warmup Training steps
Paraformer-large 4 28 5.00E-04 5000 101579
Whisper-large 1 8 5.00E-6 1000 90000
Whisper-medium 1 3 6.25e-6 1000 225000
Whisper-small 1 4 1.25e-5 1000 225000
Table 13: Hyperparameters for fine-tuning pre-trained ASR models.
Model Batch size LR Schedule Init LR Base LR Epochs
ECAPA-TDNN 128 Cyclic 5.00E-03 1.00E-08 20
ResNet-TDNN 128 Cyclic 5.00E-03 1.00E-08 20
x-vector 128 Linear 5.00E-03 1.00E-04 20
Table 14: Hyperparameters for training speaker verification models.
Model Params Batch size LR Warmup Training steps
CampNet 21.22M 16 2.00E-04 8000 200000
EditSpeech 48.15M 16 2.00E-04 8000 200000
A3T 14.86M 16 2.00E-04 8000 200000
FluentSpeech 23.86M 30 2.00E-04 8000 200000
Table 15: Hyperparameters for training speaker editing models.
Model Shanghai Guangxi Sichuan Anhui Zhejiang Jiangsu Hunan
Paraformer-fintuned 13.68 4.88 9.62 16.11 12.42 16.57 14.6
Paraformer 13.26 6.24 10.07 17.91 12.3 16.61 13.18
Brachformer 27.37 10.82 25.94 36.58 29.83 36.68 39.94
Conformer 29.11 11.91 27.11 38.01 30.17 39.44 40.65
Tranformer 42.27 26.91 41.78 52.3 46.53 51.82 54.9
Table 16: Model Performance in Southern Provinces
Model Beijing Heilongjiang Liaoning Hebei Henan
Paraformer-fintuned 7.38 5.85 27.08 7.71 20.93
Paraformer 6.8 5.38 23.85 9.41 24.19
Brachformer 23.65 21.21 55.31 19.63 40.85
Conformer 25.15 20.27 56.88 20.44 41.59
Tranformer 43.01 39.61 67.58 36.61 56.03
Table 17: Model Performance in Northern Provinces
B.1.1 Speaker Verification and Diarization
For the speaker verification experiments, we first segment the original dialogue dataset into sentence-level units. Given that the dataset includes precise timestamps for each sentence and specific audio markers for events such as overlapping speech and noise, we extract clean audio segments devoid of these special sound events and speaker overlaps. From the 101 dialogues in the dataset, we randomly select 10 dialogues, using the corresponding segmented audio at the sentence level as our test dataset. The remaining audio is split, with 10% randomly allocated for the validation set and 90% designated for the training set. The detailed information of the datasplit is shown in 11. After the datasplit, we create 20,000 carefully balanced verification pairs (50% genuine vs. 50% impostor pairs) from the test set. This trial composition ensures uniform coverage of both intra-speaker ${(spk_{i},spk_{j})}$ and inter-speaker pairs ${(spk_{i},spk_{j})}$. The detailed datasplit information is shown in table 11. For speaker diarization, the dataset split is identical to that of the speaker verification task, except that we use dialogues instead of sentences.
Split # Spk. # Utt. Dur. (hrs) Avg. (s)
Train 162 47269 29.95 2.28
Dev 20 6891 4.09 2.14
Test 20 5869 3.77 2.31
Sum 202 60029 37.81 2.27
Table 18: Summary of dataset splits, including the number of speakers (# Spk.) and utterances (# Utt.), total duration (Dur.), and average utterance length (Avg.) for speech recognition task.
B.1.2 Speech Recognition
For the elderly dataset, we initially segment the dialogue - based data into clean sentence - level data following the annotations. This process is similar to the data split in the speaker verification task described in Appendix B.1.1. Subsequently, we partition the data by speaker and randomly divide it into three subsets: the training set, the development set (dev), and the test set, with a ratio of 8:1:1. The detailed data split information is presented in Table 18.
B.1.3 Speech Editing
After converting the conversation into sentences using the method described in the speaker verification data split in Appendix B.1.1, we filter the sentences. Only those with more than four characters are included for training and testing. We then divide the data into a training-to-testing ratio of 9:1.
Split # Utt. Dur. (hrs) Avg. (s)
Train 44954 32.42 2.60
Test 4994 3.68 2.65
Sum 49948 36.10 2.60
Table 19: Summary of dataset splits, including the utterances (# Utt.), total duration (Dur.), and average utterance length (Avg.) for speech editing tasks.
B.2 Hyperparams
B.2.1 ASR Model Training from Scratch
The hyperparameters used for training ASR models from scratch are presented in Table 12. We experimented with three different encoder architectures: Transformer, Conformer, and E-Branchformer. For all models, we used an accumulation gradient of 4 and a batch size of 4. The initial learning rate was set to 1.00E-03 for Transformer and Conformer models, while it was 5.00E-04 for the E-Branchformer model. A warmup period of 5000 steps was used for all models. The models were trained for a maximum of 100 epochs (Transformer) and 60 epochs (Conformer and E-Branchformer).
B.2.2 ASR Model Fine-tuning
Table 13 outlines the hyperparameters used for fine-tuning pre-trained ASR models. Our experiments covered Paraformer-large and multiple Whisper variants (large/medium/small). The batch size varied depending on the model, ranging from 3 to 28. The learning rates were carefully chosen for each model, spanning from 5.00E-06 to 5.00E-04. A warmup period of 1000 steps was used across all fine-tuned models. The total training steps also varied significantly depending on the model, reflecting the different sizes and pre-training strategies. All the hyperparameters of whisper models are the default setting of whisper-flamingo https://github.com/roudimit/whisper-flamingo.
B.2.3 Speaker Verification Model Training
Training details for our speaker verification models – including ECAPA-TDNN, ResNet-TDNN, and x-vector architectures – are summarized in Table 14. All models used a batch size of 128. For ECAPA-TDNN and ResNet-TDNN, a cyclic learning rate schedule was employed with an initial learning rate of 5.00E-03 and a base learning rate of 1.00E-08. The x-vector model utilized a linear learning rate schedule with an initial learning rate of 5.00E-03 and a base learning rate of 1.00E-04. All models were trained for 20 epochs.
B.2.4 Speech Editing Model Training
For training our second set of Speech Editing models(CampNet, EditSpeech, A3T, and FluentSpeech), we utilized the hyperparameters summarized in Table 15. The number of trainable parameters for each model is also provided. All models, except FluentSpeech, used a batch size of 16. FluentSpeech used a batch size of 30. The initial learning rate was set to 2.00E-04 for all models. A warmup period of 8000 steps was used, and all models were trained for 200,000 steps.
B.3 Extra Experiment Analysis
This section primarily presents the detailed CER of the models trained with Wenet in the speech recognition experiments. Specifically, the specific recognition results of these models for sentences from different provinces are shown in Table 17 and Table 16. In Table 17, we can observe that the CERs of Beijing, Hebei, and Heilongjiang are relatively low, while the recognition difficulty for Liaoning and Henan increases. Similarly, in Table 16, Guangxi has the lowest recognition difficulty among all provinces. Moreover, the order of recognition difficulty, that is, the ranking of CERs, among other provinces is roughly the same. This indicates that when we break down the regions to the provincial level, the recognition difficulty varies across different provinces. Chen et al. (2024) Jia et al. (2024)